# On the Hidden Mystery of OCR in Large Multimodal Models

### v5

@@ 5 Conclusion
This paper has presented an extensive study on performance of LMM on OCR tasks, including text recognition, Scene Text-Centric, Document-Oriented VQA, KIE and HMER. Our quantitative assessment reveals that LMM can achieve promising results, especially in text recognition, even attaining SOTA performance in some datasets. However, significant gaps persist compared to domain-specific supervised methods, suggesting that specialized techniques tailored to each task are still essential, as the latter uses much less computational resources and data. However, the emergence of LMM does present promising directions for advancements in OCR techniques. It stimulates an intuitive query for potential improvements: after training on a larger amount of text-related data, can LMM surpass supervised methods? We hope this work provides readers with valuable insights into the application of LMM for future exploration in OCR technology development. Importantly, the baseline results presented in this paper can serve as a foundation for the development and evaluation of novel ideas aimed at improving text-related multimodal techniques.
@@ 6 Acknowledgements
Thanks for the help from Hongliang Li, Wenwen Yu, Yang Liu, Mingxin Huang, Dezhi Peng, Mingyu Liu and Mingrui Chen.

### v6

@@ 5 Conclusion
This paper has presented an extensive study on the performance of LMM on OCR tasks, including text recognition, Scene Text-Centric, Document-Oriented VQA, KIE, and HMER. Our quantitative assessment reveals that LMM can achieve promising results, especially in text recognition, even attaining SOTA performance in some datasets. However, significant gaps persist compared to domain-specific supervised methods, suggesting that specialized techniques tailored to each task are still essential, as the latter uses much less computational resources and data. The proposed OCRBench has served as the evaluation benchmark for the OCR capabilities of multimodal large models, driving the development of LMMs and demonstrating their immense potential in the OCR field. In the future, we hope to further explore the potential of LMMs across more scenarios, more complex tasks, and multiple languages, ultimately paving the way for more intelligent and versatile OCR solutions for large multimodal models.
@@ 6 Acknowledgements
This work was supported by National Natural Science Foundation of China (Grant Nos. 62225603 and 62226104).Thanks for the help from Hongliang Li, Yang Liu, Dezhi Peng, Mingyu Liu and Mingrui Chen.